5. Daves nominal wage is $15 per hour in 2016. year. Daves real wage in 2016 is a. $12.5, 2009 b. $12.5, 2016 c. $8, 2009 d. $8, 2016 e. $15, 2009 The CPI price index for 2016 is 120 with 2009 as the base per hour irn dollars. 6. Which one of following economists wrote The General Theory, a book that stressed the important role of the government in maintaining the stability of the macroeconomy a. Janet Yellen b. Adam Smith c. John Maynard Keynes d. Ben Bernanke e. Alan Blinder 7. The following statistics are for a nation called Wonderworld in year 1: Real GDP- $10,000,000, Population 5000, Hours of work per year 250,000 hours. From the data above, the real GDP per capita is a. $40 per hour of work; $2,000 per person. b. $ c. $2,000 per hour of work; $40 per person. d. $2,000 per person; $40 per hour of work. and labor productivity is 40 per person; $2,000 per hour of work.
Answer #1

5. a. 12.5, 2009

Real wage in a year = (Nominal wage in a year/CPI in a year) * 100

Real wage in 2016 = (15/120) * 100 = 0.125*100 = 12.5

6. c. John Maynard Keynes

The General Theory of Employment, Interest and Money of 1936 by the Economist John Maynard Keynes.

7. d. $2000 per person; $40 per hour of work.

Real GDP per capita = Real GDP/Population = 10,000,000/5000 = 2000

Labor Productivity = Real GDP /Hours of work per year = 10,000,000/250000 = 40